OpenStudy (anonymous):

64 percent Tom is riding his bicycle. He rides at a speed of 11 kilometers per hour for 17.6 kilometers. For how many hours does he ride?

OpenStudy (anonymous):

Hi! Let's start with an easier problem: if he rides at 11 kilometers per hour for 11 kilometers, for how many kilometers does he ride?

OpenStudy (anonymous):

Ooops, I meant "for how many hours does he ride?" Sorry!!

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Great! Now suppose it was for 22 kilometers?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Exactly! How about 55 kilometers?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

That's right! Now lets go back to your question: what about 17.6 kilometers?

OpenStudy (anonymous):

(This is a bit harder than 55 km, but it's the same idea. Do you need a hint?)

OpenStudy (anonymous):

1.6?

OpenStudy (anonymous):

Perfect! Well done! :)

jhonyy9 (jhonyy9):

64 percent from the first line what mean ?

OpenStudy (anonymous):

thanks

jhonyy9 (jhonyy9):

one mor easy way to solve it using the rule of 3 simple so write 11 km -------------- 1 hour 17,6 km ------------- x ------------------------------- x = 17,6 *1 / 11 = 17,6/11 = 1,6 hour x=1,6 hour hope this is understandably too